### 29 Jan 2024 v1

# High-Quality Image Restoration Following Human Instructions

Abstract
Image restoration is a fundamental problem that involves recovering a high-quality clean image from its degraded observation. All-In-One image restoration models can effectively restore images from various types and levels of degradation using degradation-specific information as prompts to guide the restoration model. In this work, we present the first approach that uses human-written instructions to guide the image restoration model. Given natural language prompts, our model can recover high-quality images from their degraded counterparts, considering multiple degradation types. Our method, InstructIR, achieves state-of-the-art results on several restoration tasks including image denoising, deraining, deblurring, dehazing, and (low-light) image enhancement. InstructIR improves +1dB over previous all-in-one restoration methods. Moreover, our dataset and results represent a novel benchmark for new research on text-guided image restoration and enhancement.
Figure 1: Given an image and a prompt for how to improve that image, our all-in-one restoration model corrects the image considering the human instruction. InstructIR, can tackle various types and levels of degradation, and it is able to generalize to restoring real photos.
1 Introduction
Images often contain unpleasant effects such as noise, motion blur, haze, and low dynamic range. Such effects are commonly known in low-level computer vision as degradations. These can result from camera limitations or challenging environmental conditions e.g. low light.
Image restoration aims to recover a high-quality image from its degraded counterpart. This is a complex inverse problem since multiple different solutions can exist for restoring any given image [20, 59, 102, 103, 16, 44].
Some methods focus on specific degradations, for instance reducing noise (denoising) [102, 103, 64], removing blur (deblurring) [58, 105], or clearing haze (dehazing) [66, 16]. Such methods are effective for their specific task, yet they do not generalize well to other types of degradation. Other approaches use a general neural network for diverse tasks [74, 95, 82, 10], yet training the neural network for each specific task independently. Since using a separate model for each possible degradation is resource-intensive, recent approaches propose All-in-One restoration models [42, 61, 60, 100]. These approaches use a single deep blind restoration model considering multiple degradation types and levels. Contemporary works such as PromptIR [61] or ProRes [49] utilize a unified model for blind image restoration using learned guidance vectors, also known as “prompt embeddings”, in contrast to raw user prompts in text form, which we use in this work.
In parallel, recent works such as InstructPix2Pix [4] show the potential of using text prompts to guide image generation and editing models. However, this method (or recent alternatives) do not tackle inverse problems. Inspired by these works, we argue that text guidance can help to guide blind restoration models better than the image-based degradation classification used in previous works [42, 100, 60]. Users generally have an idea about what has to be fixed (though they might lack domain-specific vocabulary) so we can use this information to guide the model.
Contributions
We propose the first approach that utilizes real human-written instructions to solve inverse problems and image restoration. Our comprehensive experiments demonstrate the potential of using text guidance for image restoration and enhancement by achieving state-of-the-art performance on various image restoration tasks, including image denoising, deraining, deblurring, dehazing and low-light image enhancement. Our model, InstructIR, is able to generalize to restoring images using arbitrary human-written instructions. Moreover, our single all-in-one model covers more tasks than many previous works. We show diverse restoration samples of our method in Figure 1.
2 Related Work
Image Restoration.
Recent deep learning methods [16, 64, 58, 44, 95, 74] have shown consistently better results compared to traditional techniques for blind image restoration [29, 18, 73, 35, 54, 37]. The proposed neural networks are based on convolutional neural networks (CNNs) and Transformers [76] (or related attention mechanisms).
We focus on general-purpose restoration models [44, 95, 82, 10]. For example, SwinIR [44], MAXIM [74] and Uformer [82]. These models can be trained -independently- for diverse tasks such as denoising, deraining or deblurring. Their ability to capture local and global feature interactions, and enhance them, allows the models to achieve great performance consistently across different tasks. For instance, Restormer [95] uses non-local blocks [79] to capture complex features across the image.
NAFNet [10] is an efficient alternative to complex transformer-based methods. The model uses simplified channel attention, and gating as an alternative to non-linear activations. The builing block (NAFBlock) follows a simple meta-former [92] architecture with efficient inverted residual blocks [31]. In this work, we build our InstructIR model using NAFNet as backbone, due to its efficient and simple design, and high performance in several restoration tasks.
All-in-One Image Restoration.
Single degradation (or single task) restoration methods are well-studied, however, their real-world applications are limited due to the required resources i.e. allocating different models, and select the adequate model on demand. Moreover, images rarely present a single degradation, for instance noise and blur are almost ubiquitous in any image capture.
All-in-One (also known as multi-degradation or multi-task) image restoration is emerging as a new research field in low-level computer vision [42, 61, 60, 97, 98, 49, 91, 75]. These approaches use a single deep blind restoration model to tackle different degradation types and levels.
We use as reference AirNet [42], IDR [100] and ADMS [60]. We also consider the contemporary work PromptIR [61]. The methods use different techniques to guide the blind model in the restoration process. For instance, an auxiliary model for degradation classification [42, 60], or multi-dimensional guidance vectors (also known as “prompts”) [61, 49] that help the model to discriminate the different types of degradation in the image.
Despite it is not the focus of this work, we acknowledge that real-world image super-resolution is a related problem [44, 106, 48, 12], since the models aim to solve an inverse problem considering multiple degradations (blur, noise and downsampling).
Text-guided Image Manipulation.
In the recent years, multiple methods have been proposed for text-to-image generation and text-based image editing works [4, 53, 70, 34, 30]. These models use text prompts to describe images or actions, and powerful diffusion-based models for generating the corresponding images. Our main reference is InstructPix2Pix [4], this method enables editing from instructions that tell the model what action to perform, as opposed to text labels, captions or descriptions of the input or output images. Therefore, the user can transmit what to do in natural written text, without requiring to provide further image descriptions or sample reference images.
Figure 2: We train our blind image restoration models using common image datasets, and prompts generated using GPT-4, note that this is (self-)supervised learning. At inference time, our model generalizes to human-written instructions and restores real images.
3 Image Restoration Following Instructions
We treat instruction-based image restoration as a supervised learning problem similar to previous works [4]. First, we generate over 10000 prompts using GPT-4 based on our own sample instructions. We explain the creation of the prompt dataset in Sec. 3.1. We then build a large paired training dataset of prompts and degraded/clean images. Finally, we train our InstructIR model, and we evaluate it on a wide variety of instructions including real human-written prompts. We explain our text encoder in Sec 3.2, and our complete model in Sec. 3.3.
3.1 Generating Prompts for Training
Why instructions? Inspired by InstructPix2Pix [4], we adopt human written instructions as the mechanism of control for our model. There is no need for the user to provide additional information, such as example clean images, or descriptions of the visual content. Instructions offer a clear and expressive way to interact, enabling users to pinpoint the unpleasant effects (degradations) in the images.
Handling free-form user prompts rather than fixed degradation-specific prompts increases the usability of our model for laypeople who lack domain expertise. We thus want our model to be capable of understanding diverse prompts posed by users “in-the-wild” e.g. kids, adults, or photographers. To this end, we use a large language model (i.e., GPT-4) to create diverse requests that might be asked by users for the different degradations types. We then filter those generated prompts to remove ambiguous or unclear prompts (e.g., “Make the image cleaner”, “improve this image”). Our final instructions set contains over 10000 different prompts in total, for 7 different tasks. We display some examples in Table 1. As we show in Figure 2 the prompts are sampled randomly depending on the input degradation.
Table 1: Examples of our curated GPT4-generated user prompts with varying language and domain expertise.
Degradation Prompts
Denoising Can you clean the dots from my image?
Fix the grainy parts of this photo
Remove the noise from my picture
Deblurring Can you reduce the movement in the image?
My picture’s not sharp, fix it
Deblur my picture, it’s too fuzzy
Dehazing Can you make this picture clearer?
Help, my picture is all cloudy
Remove the fog from my photo
Deraining I want my photo to be clear, not rainy
Clear the rain from my picture
Remove the raindrops from my photo
Super-Res. Make my photo bigger and better
Add details to this image
Increase the resolution of this photo
Low-light The photo is too dark, improve exposure
Increase the illumination in this shot
My shot has very low dynamic range
Enhancement Make it pop!
Adjust the color balance for a natural look
Apply a cinematic color grade to the photo
3.2 Text Encoder
The Choice of the Text Encoder.
A text encoder maps the user prompt to a fixed-size vector representation (a text embedding). The related methods for text-based image generation [67] and manipulation [4, 3] often use the text encoder of a CLIP model [62] to encode user prompts as CLIP excels in visual prompts. However, user prompts for degradation contain, in general, little to no visual content (e.g. the use describes the degradation, not the image itself), therefore, the large CLIP encoders (with over 60 million parameters) are not suitable – especially if we require efficiency.
We opt, instead, to use a pure text-based sentence encoder [63], that is, a model trained to encode sentences in a semantically meaningful embedding space. Sentence encoders –pre-trained with millions of examples– are compact and fast in comparison to CLIP, while being able to encode the semantics of diverse user prompts. For instance, we use the BGE-micro-v2 sentence transformer.
Fine-tuning the Text Encoder.
We want to adapt the text encoder $\mathrm{E}$ for the restoration task to better encode the required information for the restoration model. Training the full text encoder is likely to lead to overfitting on our small training set and lead to loss of generalization. Instead, we freeze the text encoder and train a projection head on top:
$\mathbf{e}=\mathrm{norm}(\mathbf{W}\cdot\mathrm{E}(t))$ (1)
where $t$ is the text, $\mathrm{E}(t)$ represents the raw text embedding, $\mathbf{W}\in\mathbb{R}^{d_{t}\times d_{v}}$ is a learned projection from the text dimension ($d_{t}$) to the input dimension for the restoration model ($d_{v}$), and $\mathrm{norm}$ is the l2-norm.
Figure 3 shows that while the text encoder is capable out-of-the-box to cluster instructions to some extent (Figure 2(a)), our trained projection yields greatly improved clusters (Figure 2(b)). We distinguish clearly the clusters for deraining, denoising, dehazing, deblurring, and low-light image enhancement. The instructions for such tasks or degradations are very characteristic. Furthermore, we can appreciate that “super-res” and “enhancement” tasks are quite spread and between the previous ones, which matches the language logic. For instance “add details to this image” could be used for enhancement, deblurring or denosising. In our experiments, $d_{t}\!=\!384$, $d_{v}\!=\!256$ and $\mathbf{W}$ is a linear layer.
(a) t-SNE of embeddings before training i.e. frozen text encoder
(b) t-SNE of embeddings after training our learned projection
Figure 3: We show t-SNE plots of the text embeddings before/after training InstructIR. Each dot represents a human instruction.
Intent Classification Loss.
We propose a guidance loss on the text embedding $\mathbf{e}$ to improve training and interpretability. Using the degradation types as targets, we train a simple classification head $\mathcal{C}$ such that $\mathbf{c}=\mathcal{C}(\mathbf{e})$, where $\mathbf{c}\in\mathrm{R}^{D}$, being $D$ is the number of degradations. In our experiments $D=7$ and the classification head $\mathcal{C}$ is a simple two-layers MLP. Thus, we only need to train a projection layer and a simple MLP to capture the natural language knowledge. This allows the text model to learn meaningful embeddings as we can appreciate in Figure 3, not just guidance vectors for the main image processing model.
We find that the model is able to classify accurately (i.e. over 95% accuracy) the underlying degradation in the user’s prompt after a few epochs.
3.3 InstructIR
Figure 4: Instruction Condition Block (ICB) using an approximation of task routing [71] for many-tasks learning. See Eq. 2.
Our method InstructIR consists of an image model and a text encoder. We introduced our text encoder in Sec. 3.2. We use NAFNet [10] as the image model, an efficient image restoration model that follows a U-Net architecture [68]. To successfully learn multiple tasks using a single model, we use task routing techniques. Our framework for training and evaluating the model is illustrated in Figure 2.
Text Guidance.
The key aspect of InstructIR is the integration of the encoded instruction as a mechanism of control for the image model. Inspired in task routing for many-task learning [69, 71, 14], we propose an “Instruction Condition Block” (ICB) to enable task-specific transformations within the model. Conventional task routing [71] applies task-specific binary masks to the channel features. Since our model does not know a-priori the degradation, we cannot use this technique directly.
Considering the image features $\mathcal{F}$, and the encoded instruction $\mathbf{e}$, we apply task routing as follows:
$\mathcal{F^{\prime}}_{c}=\mathrm{Block}(\mathcal{F}_{c}\odot\mathbf{m}_{c})+\mathcal{F}_{c}$ (2)
where the mask $\mathbf{m}_{c}=\sigma(\mathbf{W_{c}}\cdot\mathbf{e})$ is produced using a linear layer -activated using the Sigmoid function- to produce a set of weights depending on the text embedding $\mathbf{e}$. Thus, we obtain a $c$-dimensional per-channel (soft-)binary mask $\mathbf{m}_{c}$. As [71], task routing is applied as the channel-wise multiplication $\odot$ for masking features depending on the task. The conditioned features are further enhanced using a NAFBlock [10] ($\mathrm{Block}$). We illustrate our task-routing ICB block in Figure 4. We use “regular” NAFBlocks [10], followed by ICBs to condition the features, at both encoder and decoder blocks. The formulation is $F^{l+1}\!=\!\mathrm{ICB}(\mathrm{Block}(F^{l}))$ where $l$ is the layer. Although we do not condition explicitly the filters of the neural network, as in [71], the mask allows the model to select the most relevant channels depending on the image information and the instruction. Note that this formulation enables differentiable feature masking, and certain interpretability i.e. the features with high weights contribute the most to the restoration process. Indirectly, this also enforces to learn diverse filters and reduce sparsity [71, 14].
Is InstructIR a blind restoration model?
The model does not use explicit information about the degradation in the image e.g. noise profiles, blur kernels, or PSFs. Since our model infers the task (degradation) given the image and the instruction, we consider InstructIR a blind image restoration model. Similarly to previous works that use auxiliary image-based degradation classification [60, 42].
4 Experimental Results
We provide extensive qualitative results using benchmark images in Figures 1718 and 19. We also evaluate our model on 9 well-known benchmarks for different image restoration tasks: image denoising, deblurring, deraining, dehazing, and image enhancement. We present extensive quantitative results in Table 2. Our single model successfully restores images considering different degradation types and levels. We provide additional results and ablation studies in the supplementary material.
4.1 Implementation Details.
Our InstructIR model is end-to-end trainable. The image model does not require pre-training, yet we use a pre-trained sentence encoder as language model.
Text Encoder.
As we discussed in Sec. 3.2, we only need to train the text embedding projection and classification head ($\approx\!100K$ parameters). We initialize the text encoder with BGE-micro-v2  [^1] https://huggingface.co/TaylorAI/bge-micro-v2, a distilled version of BGE-small-en [85]. The BGE encoders are BERT-like encoders [13] pre-trained on large amounts of supervised and unsupervised data for general-purpose sentence encoding. The BGE-micro model is a 3-layer encoder with 17.3 million parameters, which we freeze during training. We also explore all-MiniLM-L6-v2 and CLIP encoders, however, we concluded that small models prevent overfitting and provide the best performance while being fast. We provide the ablation study comparing the three text encoders in the supplementary material.
Image Model.
We use NAFNet [10] as image model. The architecture consists of a 4-level encoder-decoder, with varying numbers of blocks at each level, specifically [2, 2, 4, 8] for the encoder, and [2, 2, 2, 2] for the decoder, from the level-1 to level-4 respectively. Between the encoder and decoder we use 4 middle blocks to enhance further the features. The decoder implements addition instead of concatenation for the skip connections.
We use the Instruction Condition Block (ICB) for task-routing [71] only in the encoder and decoder.
The model is optimized using the $\mathcal{L}_{1}$ loss between the ground-truth clean image and the restored one. Additionally we use the cross-entropy loss $\mathcal{L}_{ce}$ for the intent classification head of the text encoder. We train use a batch size of 32 and AdamW [36] optimizer with learning rate $5e^{-4}$ for 500 epochs (approximately 1 day using a single NVIDIA A100). We also use cosine annealing learning rate decay. During training, we utilize cropped patches of size $256\times 256$ as input, and we use random horizontal and vertical flips as augmentations. Since our model uses as input instruction-image pairs, given an image, and knowing its degradation, we randomly sample instructions from our prompt dataset ($>\!10$K samples). Our image model has only 16M parameters, and the learned text projection is just $100$k parameters (the language model is 17M parameters), thus, our model can be trained easily on standard GPUs such as NVIDIA RTX 2080Ti or 3090Ti in a couple of days. Furthermore, the inference process also fits in low-computation budgets.
4.2 Datasets and Benchmarks
Following previous works [42, 100, 61], we prepare the datasets for different restoration tasks.
Image denoising.
We use a combination of BSD400 [2] and WED [50] datasets for training. This combined training set contains $\approx\!5000$ images. Using as reference the clean images in the dataset, we generate the noisy images by adding Gaussian noise with different noise levels $\sigma\in\{15,25,50\}$. We test the models on the well-known BSD68 [52] and Urban100 [32] datasets.
Image deraining.
We use the Rain100L [88] dataset, which consists of 200 clean-rainy image pairs for training, and 100 pairs for testing.
Image dehazing.
We utilize the Reside (outdoor) SOTS [41] dataset, which contains $\approx\!72$K training images. However, many images are low-quality and unrealistic, thus, we filtered the dataset and selected a random set of 2000 images – also to avoid imbalance w.r.t the other tasks. We use the standard outdoor testset of 500 images.
Image deblurring.
We use the GoPro dataset for motion deblurring [57] which consist of 2103 images for training, and $1$ for testing.
Low-light Image Enhancement.
We use the LOL [83] dataset (v1), and we adopt its official split of 485 training images, and 15 testing images.
Image Enhancement.
Extending previous works, we also study photo-realistic image enhancement using the MIT5K dataset [5]. We use 1000 images for training, and the standard split of 500 images for testing (as in [74]).
Table 2: Quantitative results on five restoration tasks (5D) with state-of-the-arts general image restoration and all-in-one methods. We highlight the reference model without text, the best results, and the second best results. We also present the ablation study of our multi-task variants (from 5 to 7 tasks — 5D, 6D, 7D). This table is based on Zhang et al. IDR [100] (CVPR ’23).
Deraining Dehazing Denoising Deblurring Low-light Enh.
Methods Rain100L [88] SOTS [41] BSD68 [52] GoPro [57] LOL [83] Average Params
PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ PSNR↑ SSIM↑ (M)
HINet [9] 35.67 0.969 24.74 0.937 31.00 0.881 26.12 0.788 19.47 0.800 27.40 0.875 88.67
DGUNet [56] 36.62 0.971 24.78 0.940 31.10 0.883 27.25 0.837 21.87 0.823 28.32 0.891 17.33
MIRNetV2 [93] 33.89 0.954 24.03 0.927 30.97 0.881 26.30 0.799 21.52 0.815 27.34 0.875 5.86
SwinIR [44] 30.78 0.923 21.50 0.891 30.59 0.868 24.52 0.773 17.81 0.723 25.04 0.835 0.91
Restormer [95] 34.81 0.962 24.09 0.927 31.49 0.884 27.22 0.829 20.41 0.806 27.60 0.881 26.13
NAFNet [10] 35.56 0.967 25.23 0.939 31.02 0.883 26.53 0.808 20.49 0.809 27.76 0.881 17.11
DL [21] 21.96 0.762 20.54 0.826 23.09 0.745 19.86 0.672 19.83 0.712 21.05 0.743 2.09
Transweather [75] 29.43 0.905 21.32 0.885 29.00 0.841 25.12 0.757 21.21 0.792 25.22 0.836 37.93
TAPE [45] 29.67 0.904 22.16 0.861 30.18 0.855 24.47 0.763 18.97 0.621 25.09 0.801 1.07
AirNet [42] 32.98 0.951 21.04 0.884 30.91 0.882 24.35 0.781 18.18 0.735 25.49 0.846 8.93
IDR [100] 35.63 0.965 25.24 0.943 31.60 0.887 27.87 0.846 21.34 0.826 28.34 0.893 15.34
InstructIR-5D 36.84 0.973 27.10 0.956 31.40 0.887 29.40 0.886 23.00 0.836 29.55 0.907 15.8
InstructIR-6D 36.80 0.973 27.00 0.951 31.39 0.888 29.73 0.892 22.83 0.836 29.55 0.908 15.8
InstructIR-7D 36.75 0.972 26.90 0.952 31.37 0.887 29.70 0.892 22.81 0.836 29.50 0.907 15.8
Finally, as previous works [42, 100, 61], we combine all the aforementioned training datasets, and we train our unified model for all-in-one restoration.
4.3 Multiple Degradation Results
We define two initial setups for multi-task restoration:
  • 3D for three-degradation models such as AirNet [42], these tackle image denoising, dehazing and deraining.
  • 5D for five-degradation models, considering image denoising, deblurring, dehazing, deraining and low-light image enhancement as in [100].
In Table 2, we show the performance of 5D models. Following Zhang et al. [100], we compare InstructIR with several state-of-the-art methods for general image restoration [95, 10, 9, 44, 93], and all-in-one image restoration methods [100, 42, 75, 21, 45]. We can observe that our simple image model (just 16M parameters) can tackle successfully at least five different tasks thanks to the instruction-based guidance, and achieves the most competitive results.
In Table 4 we can appreciate a similar behaviour, when the number of tasks is just three (3D), our model improves further in terms of reconstruction performance.
Based on these results, we pose the following question: How many tasks can we tackle using a single model without losing too much performance? To answer this, we propose the 6D and 7D variants. For the 6D variant, we fine-tune the original 5D to consider also super-resolution as sixth task. Finally, our 7D model includes all previous tasks, and additionally image enhancement (MIT5K photo retouching). We show the performance of these two variants in Table 2.
Table 3: Ablation study on the sensitivity of instructions. We report PSNR/SSIM metrics for each task using our 5D base model. We repeat the evaluation on each testset 10 times, each time we sample different prompts for each image, and we report the average results. The “Real Users $\dagger$” in this study are amateur photographers, thus, the instructions were very precise.
Language Level Deraining Denoising Deblurring LOL
Basic & Precise 36.84/0.973 31.40/0.887 29.47/0.887 23.00/0.836
Basic & Ambiguous 36.24/0.970 31.35/0.887 29.21/0.885 21.85/0.827
Real Users $\dagger$ 36.84/0.973 31.40/0.887 29.47/0.887 23.00/0.836
Test Instructions.
InstructIR requires as input the degraded image and the human-written instruction. Therefore, we also prepare a testset of prompts i.e. instruction-image test pairs. The performance of InstructIR depends on the ambiguity and precision of the instruction. We provide the ablation study in Table 3. InstructIR is quite robust to more/less detailed instructions. However, it is still limited with ambiguous instructions such as “enhance this image”. We show diverse instructions in the following Figures.
Table 4: Comparisons of all-in-one restoration models for three restoration tasks (3D). We also show an ablation study for image denoising -the fundamental inverse problem- considering different noise levels. We report PSNR/SSIM metrics. Table based on [61].
Methods Dehazing Deraining Denoising ablation study (BSD68 [52]) Average
SOTS [41] Rain100L [21] $\sigma=15$ $\sigma=25$ $\sigma=50$
BRDNet [72] 23.23/0.895 27.42/0.895 32.26/0.898 29.76/0.836 26.34/0.836 27.80/0.843
LPNet [25] 20.84/0.828 24.88/0.784 26.47/0.778 24.77/0.748 21.26/0.552 23.64/0.738
FDGAN [19] 24.71/0.924 29.89/0.933 30.25/0.910 28.81/0.868 26.43/0.776 28.02/0.883
MPRNet [94] 25.28/0.954 33.57/0.954 33.54/0.927 30.89/0.880 27.56/0.779 30.17/0.899
DL[21] 26.92/0.391 32.62/0.931 33.05/0.914 30.41/0.861 26.90/0.740 29.98/0.875
AirNet [42] 27.94/0.962 34.90/0.967 33.92/0.933 31.26/0.888 28.00/0.797 31.20/0.910
PromptIR [61] 30.58/0.974 36.37/0.972 33.98/0.933 31.31/0.888 28.06/0.799 32.06/0.913
InstructIR-3D 30.22/0.959 37.98/0.978 34.15/0.933 31.52/0.890 28.30/0.804 32.43/0.913
InstructIR-5D 27.10/0.956 36.84/0.973 34.00/0.931 31.40/0.887 28.15/0.798 31.50/0.909
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rain, Blur and Noise “Correct the noise” “Remove the rain” “Increase resolution” “Enhance the photo”
Figure 5: Selective task. InstructIR can remove particular degradations or perform different transformations depending on the human instructions. This is a novel feature in image restoration, and it is possible thanks to the novel integration of textual descriptions.
5 Multi-Task Ablation Study
Table 5: Image Enhancement performance on MIT5K [5, 96].
Method PSNR $\uparrow$ SSIM $\uparrow$ $\Delta\!E_{ab}~{}\downarrow$
UPE [77] 21.88 0.853 10.80
DPE [26] 23.75 0.908 9.34
HDRNet [11] 24.32 0.912 8.49
3DLUT [96] 25.21 0.922 7.61
InstructIR-7D 24.65 0.900 8.20
Table 6: Summary ablation study on the multi-task variants of InstructIR that tackle from 3 to 7 tasks. We report PSNR/SSIM.
Tasks Rain Noise ($\sigma 15$) Blur LOL
3D 37.98/0.978 31.52/0.890 - -
5D 36.84/0.973 31.40/0.887 29.40/0.886 23.00/0.836
6D 36.80 0.973 31.39 0.888 29.73/0.892 22.83 0.836
7D 36.75 0.972 31.37 0.887 29.70/0.892 22.81 0.836
How does 6D work?
Besides the 5 basic tasks -as previous works-, we include single image super-resolution (SISR). For this, we include as training data the DIV2K [1]. Since our model does not perform upsampling, we use the Bicubic degradation model [1, 15] for generating the low-resolution images (LR), and the upsampled versions (HR) that are fed into our model to enhance them. Adding this extra task increases the performance on deblurring –a related degradation–, without harming notably the performance on the other tasks. However, the performance on SR benchmarks is far from classical super-resolution methods [1, 44].
How does 7D work?
Finally, if we add image enhancement –a task not related to the previous ones i.e. inverse problems– the performance on the restoration tasks decays slightly. However, the model still achieves state-of-the-art results. Moreover, as we show in Table 5, the performance on this task using the MIT5K [5] Dataset is notable, while keeping the performance on the other tasks. We achieve similar performance to classical task-specific methods.
We summarize the multi-task ablation study in Table 6. Our model can tackle multiple tasks without losing performance notably thanks to the instruction-based task routing.
Refer to caption Refer to caption Refer to caption Refer to caption
Input “Clean up my image, “Get rid of the grain “Remove the strange spots
it’s too fuzzy.” in my photo” on my photo”
Refer to caption Refer to caption Refer to caption Refer to caption
“Retouch this image and “Reduce the motion “Please get rid of “Reduce the fog in
improve colors” in this shot” the raindrops” this landmark”
Figure 6: Instruction-based Image Restoration. InstructIR understands a wide a range of instructions for a given task (first row). Given an adversarial instruction (second row), the model performs an identity –we did not enforce this during training–. Images from BSD68 [52].
Refer to caption Refer to caption Refer to caption
Input (1)“Clear the rain from my picture” $\longrightarrow$ (2)“Make this photo look breathtaking”
Refer to caption Refer to caption Refer to caption
(1) “Retouch it as a photographer” $\longrightarrow$ (2) “Can you remove the raindrops?” $\longrightarrow$ (3) “Increase the resolution and details”
Refer to caption Refer to caption Refer to caption
Input (1)“My image is too dark, can you fix it?” (2)“Apply tone-mapping to the photo”
Figure 7: Multiple Real Instructions. We can prompt multiple instructions (in sequence) to restore and enhance the images. This provides additional control. We show two examples of multiple instructions applied to the “Input” image -from left to right-.
Comparison with Task-specific Methods
Our main goal is to design a powerful all-in-one model, thus, InstructIR was not designed to be trained for a particular degradation. Nevertheless, we also compare InstructIR with task-specific methods i.e. models tailored and trained for specific tasks.
We compare with task-specific methods for image enhancement in Table 5, and for low-light in image enhancement in 7. We provide extensive comparisons for image denoising in Table 8. Also, in Table 9 we show comparisons with classical methods for deblurring and dehazing. Our multi-task method is better than most task-specific methods, yet it is still not better than SOTA.
6 On the Effectiveness of Instructions
Thanks to our integration of human instructions, users can control how to enhance the images. We show an example in Figure 5, where the input image has three different degradations, and we aim to focus on a particular one. Although these results do not offer great reconstruction, we believe it is a promising direction that illustrates the effectiveness of instruction guidance for image restoration and enhancement. We provide more results in Figures 6 and 7, where we show the potential of our method to restore and enhance images in a controllable manner.
Table 7: Quantitative comparisons with state-of-the-art methods on the LOL dataset [83] (low-light enhancement). Table based on [81].
Method LPNet [43] URetinex -Net[84] DeepLPF [55] SCI [51] LIME [27] MF [23] NPE [78] SRIE [24] SDD [28] CDEF [40] InstructIR Ours
PSNR $\uparrow$ 21.46 21.32 15.28 15.80 16.76 16.96 16.96 11.86 13.34 16.33 22.83
SSIM $\uparrow$ 0.802 0.835 0.473 0.527 0.444 0.505 0.481 0.493 0.635 0.583 0.836
Method DRBN [89] KinD [107] RUAS [46] FIDE [86] EG [33] MS-RDN [90] Retinex -Net[83] MIRNet [93] IPT [8] Uformer [82] IAGC [81]
PSNR $\uparrow$ 20.13 20.87 18.23 18.27 17.48 17.20 16.77 24.14 16.27 16.36 24.53
SSIM $\uparrow$ 0.830 0.800 0.720 0.665 0.650 0.640 0.560 0.830 0.504 0.507 0.842
Table 8: Comparison with general restoration and all-in-one methods (*) at image denoising. We report PSNR on benchmark datasets considering different $\sigma$ noise levels. Table based on [100].
CBSD68 [52] Urban100 [32] Kodak24 [22]
Method 15 25 50 15 25 50 15 25 50
IRCNN [103] 33.86 31.16 27.86 33.78 31.20 27.70 34.69 32.18 28.93
FFDNet [104] 33.87 31.21 27.96 33.83 31.40 28.05 34.63 32.13 28.98
DnCNN [101] 33.90 31.24 27.95 32.98 30.81 27.59 34.60 32.14 28.95
NAFNet [10] 33.67 31.02 27.73 33.14 30.64 27.20 34.27 31.80 28.62
HINet [9] 33.72 31.00 27.63 33.49 30.94 27.32 34.38 31.84 28.52
DGUNet [56] 33.85 31.10 27.92 33.67 31.27 27.94 34.56 32.10 28.91
MIRNetV2 [93] 33.66 30.97 27.66 33.30 30.75 27.22 34.29 31.81 28.55
SwinIR [44] 33.31 30.59 27.13 32.79 30.18 26.52 33.89 31.32 27.93
Restormer [95] 34.03 31.49 28.11 33.72 31.26 28.03 34.78 32.37 29.08
* DL [21] 23.16 23.09 22.09 21.10 21.28 20.42 22.63 22.66 21.95
* T.weather [75] 31.16 29.00 26.08 29.64 27.97 26.08 31.67 29.64 26.74
* TAPE [45] 32.86 30.18 26.63 32.19 29.65 25.87 33.24 30.70 27.19
* AirNet [42] 33.49 30.91 27.66 33.16 30.83 27.45 34.14 31.74 28.59
* IDR [100] 34.11 31.60 28.14 33.82 31.29 28.07 34.78 32.42 29.13
* InstructIR-5D 34.00 31.40 28.15 33.77 31.40 28.13 34.70 32.26 29.16
* InstructIR-3D 34.15 31.52 28.30 34.12 31.80 28.63 34.92 32.50 29.40
This implies an advancement w.r.t classical (deterministic) image restoration methods. Classical deep restoration methods lead to a unique result, thus, they do not allow to control how the image is processed. We also compare InstructIR with InstructPix2Pix [4] in Figure 8.
Qualitative Results.
We provide diverse qualitative results for several tasks. In Figure 9, we show results on the LOL [83] dataset. In Figure 10, we compare methods on the motion deblurring task using the GoPro [57] dataset. In Figure 11, we compare with different methods for the dehazing task on SOTS (outdoor) [41]. In Figure 12, we compare with image restoration methods for deraining on Rain100L [21]. Finally, we show denoising results in Figure 13. In this qualitative analysis, we use our single InstructIR-5D model to restore all the images.
Table 9: Deblurring and Dehazing comparisons. We compare with task-specific classical methods on benchmark datasets.
Deblurring GoPro [57] Dehazing SOTS [41]
Method PSNR/SSIM Method PSNR/SSIM
Xu et al. [87] 21.00/0.741 DehazeNet [6] 22.46/0.851
DeblurGAN [38] 28.70/0.858 GFN [65] 21.55/0.844
Nah et al. [57] 29.08/0.914 GCANet [7] 19.98/0.704
RNN [99] 29.19/0.931 MSBDN [17] 23.36/0.875
DeblurGAN-v2 [39] 29.55/0.934 DuRN  [47] 24.47/0.839
InstructIR-5D 29.40/0.886 InstructIR-5D 27.10/0.956
InstructIR-6D 29.73/0.892 InstructIR-3D 30.22/0.959
Discussion on Instruction-based Restoration
In Figure 8 we compare with InstructPix2Pix [4]. Our method is notably superior in terms of efficiency, fidelity and quality. We can conclude that diffusion-based methods [67, 53, 4] for image manipulation require complex “tuning” of several (hyper-)parameters, and strong regularization to enforce fidelity and reduce hallucinations. InstructPix2Pix [4] cannot solve inverse problems directly –although it has a good prior for solving Inpainting–, which indicates that such model require restoration-specific training (or fine-tuning).
Limitations
Our method achieves state-of-the-art results in five tasks, proving the potential of using instructions to guide deep blind restoration models. However, we acknowledge certain limitations. First, in comparison to diffusion-based restoration methods, our current approach would not produce better results attending to perceptual quality. Second, our model struggles to process images with more than one degradation (i.e. real-world images), yet this is a common limitation among the related restoration methods. Third, as previous all-in-one methods, our model only works with in-distribution degradations, thus it will not work on unseen artifacts. Nevertheless, these limitations can be surpassed with more realistic training data.
7 Conclusion
We present the first approach that uses human-written instructions to guide the image restoration models. Given natural language prompts, our model can recover high-quality images from their degraded counterparts, considering multiple degradation types. InstructIR achieves state-of-the-art results on several restoration tasks, demonstrating the power of instruction guidance. These results represent a novel benchmark for text-guided image restoration.n
Acknowledgments This work was partly supported by the The Humboldt Foundation (AvH). Marcos Conde is also supported by Sony Interactive Entertainment, FTG.
Instruction: “Reduce the noise in this photo” – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Remove the tiny dots in this image” – Basic & Ambiguous
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Improve the quality of this image” – Real user (Ambiguous)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “restore this photo, add details” – Real user (Precise)
Refer to caption Refer to caption Refer to caption Refer to caption
Instruction: “Enhance this photo like a photographer” – Basic & Precise
Refer to caption Refer to caption Refer to caption Refer to caption
Input InstructIR (ours) InstructPix2Pix  $S_{I}\!=\!5$ InstructPix2Pix  $S_{I}\!=\!7$
Figure 8: Comparison with InstructPix2Pix [4] for instruction-based restoration using the prompt. Images from the RealSRSet [80, 44]. We use our 7D variant. We run InstructPix2Pix [4] using two configurations where we vary the weight of the image component hoping to improve fidelity: $S_{I}\!=\!5$ and $S_{I}\!=\!7$ (also known as Image CFG), this parameters helps to enforce fidelity and reduce hallucinations.
Figure 9: Low-light Image Enhancement Results. We compare with other methods on LOL [83] (748.png).
Figure 10: Image Deblurring Results. Comparison with other methods on the GoPro [57] dataset (GOPR0854-11-00-000001.png).
Figure 11: Image Dehazing Results. Comparison with other methods on SOTS [41] outdoor (0150.jpg).
Figure 12: Image Deraining Results on Rain100L [21] (035.png).
Figure 13: Image Denoising Results on BSD68 [52] (0060.png).
[…]
Supplementary Material
We define our loss functions in the paper Sec. 4.1. Our training loss function is $\mathcal{L}=\mathcal{L}_{1}+\mathcal{L}_{ce}$, which includes the loss function of the image model ($\mathcal{L}_{1}$), and the loss function for intent (task/degradation) classification ($\mathcal{L}_{ce}$) given the prompt embedding. We provide the loss evolution plots in Figures 14 and 15. In particular, in Figure 15 we can observe how the intent classification loss (i.e. predicting the task (or degradation) given the prompt), tends to 0 very fast, indicating that our language model component can infer easily the task given the instruction.
Additionally, we study three different text (sentence) encoders: (i) BGE-micro-v2  [^2] https://huggingface.co/TaylorAI/bge-micro-v2, (ii) all-MiniLM-L6-v2  [^3] https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2, (iii) CLIP text encoder (OpenAI CLIP ViT B-16). Note that these are always frozen. We use pre-trained weights from HuggingFace.
In Table 10 we show the ablation study. There is no significant difference between the text encoders. This is related to the previous results (Fig. 15), any text encoder with enough complexity can infer the task from the prompt. Therefore, we use BGE-micro-v2, as it is just 17M parameters in comparison to the others (40-60M parameters). Note that for this ablation study, we keep fixed the image model (16M), and we only change the language model.
Text Discussion
We shall ask, do the text encoders perform great because the language and instructions are too simple?
We believe our instructions cover a wide range of expressions (technical, common language, ambiguous, etc). The language model works properly on real-world instructions. Therefore, we believe the language for this specific task is self-constrained, and easier to understand and to model in comparison to other “open” tasks such as image generation.
Model Design
Based on our experiments, given a trained text-guided image model (e.g. based on NAFNet [10]), we can switch language models without performance loss.
Comparison of NAFNet with and without using text (i.e. image only): The reader can find the comparison in the main paper Table 2, please read the highlighted caption.
How the 6D variant does Super-Resolution?. We degraded the input images by downsampling and re-upsampling using Bicubic interpolation. Given a LR image, we updample it using Bicubic, then InstructIR can recover some details.
Table 10: Ablation study on the text encoders. We report PSNR/SSIM metrics for each task using our 5D base model. We use the same fixed image model (based on NAFNet [10]).
Encoder Deraining Denoising Deblurring LOL
BGE-micro 36.84/0.973 31.40/0.887 29.40/0.886 23.00/0.836
ALL-MINILM 36.82/0.972 31.39/0.887 29.40/0.886 22.98/0.836
CLIP 36.83/0.973 31.39/0.887 29.40/0.886 22.95/0.834
Figure 14: Image Restoration Loss ($\mathcal{L}_{1}$) computed between the restored image $\hat{x}$ (model’s output) and the reference image $x$.
Figure 15: Intent Classification Loss from the instructions. Product of our simple MLP classification head using $\mathbf{e}$. When $\mathcal{L}_{ce}\!\to\!0$ the model uses the learned (optimized) prompt embeddings, and it is optimized mainly based on the image regression loss ($\mathcal{L}_{1}$).
Real-World Generalization.
We evaluate InstructIR as previous works [61, 42, 100]. Also, we find the same limitations as such methods when we process real-world images. Evaluating the model on (multiple) real-world degradations is a future task.
Contemporary Works and Reproducibility.
Note that PromptIR, ProRes [49] and Amirnet [98] are contemporary works (presented or published by Dec 2023). We compare mainly with AirNet [42] since the model and results are open-source, and it is a reference all-in-one method. To the best of our knowledge, IDR [100] and ADMS [60] do not provide open-source code, models or results, thus we cannot compare with them qualitatively.
Additional Visual Results
We present diverse qualitative samples in Figures 1718, and 19. Our method produces high-quality results given images with any of the studied degradations. In most cases the results are better than the reference all-in-one model AirNet [42]. Download all the test results at https://github.com/mv-lab/InstructIR.
Refer to caption Refer to caption Refer to caption Refer to caption
Input (RealSRSet) InstructIR InstructPix2Pix  #1 InstructPix2Pix  #2
Figure 16: Comparison with InstructPix2Pix [4] for instruction-based restoration using the prompt “Remove the noise in this photo”.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input AirNet [42] PromptIR [61] InstructIR Reference
Figure 17: Denoising results for all-in-one methods. Images from BSD68 [52] with noise level $\sigma=25$.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input AirNet [42] PromptIR [61] InstructIR Reference
Figure 18: Image deraining comparisons for all-in-one methods on images from the Rain100L dataset [21].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Input AirNet [42] PromptIR [61] InstructIR Reference
Figure 19: Dehazing comparisons for all-in-one methods on images from the SOTS outdoor dataset [41].